OpenStudy (anonymous):

What is the slope of the line that passes through the points (3, –4) and (–7, 1)? A.-2 B.-1/2 C.1/2 D.2 @jelmar @hartnn @IMstuck

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

slope, m=dy/dx, the change in the dependent variable over the change in the independent variable. so m = y2-y1/x2-x1 = (-4-1)/3--7=-5/10 =-1/2

OpenStudy (imstuck):

slope formula is \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]Just fill in the y's from both coordinates, and the x's from both coordinates and subtract them. If you are struggling with it, please ask!

OpenStudy (anonymous):

its b?

OpenStudy (anonymous):

@IMStuck

OpenStudy (muzzack):

do the formula

OpenStudy (muzzack):

\[\frac{ 1-(-4) }{-7-3}\]

OpenStudy (muzzack):

this would be B